OpenStudy (anonymous):

find the magnitude of the vector <-2,-4> Please help :)

jimthompson5910 (jim_thompson5910):

for any vector v = <a,b> the magnitude, call it m, is m = |v| = sqrt(a^2 + b^2)

OpenStudy (anonymous):

so.. sqrt(20) -thats all?

jimthompson5910 (jim_thompson5910):

you can simplify that

jimthompson5910 (jim_thompson5910):

hint: 20 = 4*5

jimthompson5910 (jim_thompson5910):

another hint: sqrt(x*y) = sqrt(x) * sqrt(y)

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

2sqrt5

jimthompson5910 (jim_thompson5910):

perfect